The height of a triangle is 3 times as long as it's base. If the base is increased by 6cm. and the height is decreased by 5cm. Then the area of the triangle would be 200cm2. What is the height and base of the Original Triangle?
Answer by lwsshak3(11628)

The height of a triangle is 3 times as long as it's base. If the base is increased by 6cm. and the height is decreased by 5cm. Then the area of the triangle would be 200cm2. What is the height and base of the Original Triangle?
**
let x=base of original triangle
3x=height of original triangle
x+6 =base of new triangle
3x-5=height of new triangle
area=1/2 base*height
200=1/2*(x+6)(3x-5)
(x+6)(3x-5)=400
x^2+13x-30=400
3x^2+13x-430=0
Solve by quadratic formula:
x=-43/3 (reject, x>0)
or x=10
3x=30z
ans:
base of original triangle=10 cm
height of original triangle=30 cm